OpenStudy (anonymous):

The graph of y ≤ -x^2 - 6x - 5 contains which of the following points?

OpenStudy (texaschic101):

is this multiple choice ? Please list the answer choices, if it is.

OpenStudy (anonymous):

A. (-3, 5) B. (-3, 4) C. (1, 6) D. (2, 0)

OpenStudy (texaschic101):

now we are gonna sub and find the right answer y <= x^2 - 6x - 5 (-3,5) x = -3 and y = 5 now sub 5 <= -3^2 - 6(-3) - 5 5 <= 9 + 18 - 5 5 <= 27 - 5 5 <= 22 (correct) Can there be more then one answer ? If not, this is your answer

OpenStudy (anonymous):

thank you, it was easier than it looked

OpenStudy (texaschic101):

wait a second...I think I messed up...hold on

OpenStudy (texaschic101):

y <= -x^2 - 6x - 5 (-3,5) 5 <= -(-3^2) - 6(-3) - 5 5 <= -9 + 18 - 5 5 <= -14 + 18 5 <= 4 (incorrect) (-3,4) 4 <= -(-3) - 6(-3) - 5 4 <= -9 + 18 - 5 4 <= -14 + 18 4 <= 4 (correct) (1,6) 6 <= -1^2 - 6(1) - 5 6 <= 1 - 6 - 5 6 <= 1 - 11 6 <= -10 (incorrect) (2,0) 0 <= -2^2 - 6(2) - 5 0 <= 4 - 12 - 5 0 <= 4 - 17 0 <= - 13 (incorrect) sorry about that...messed up with the negative sign. Your answer is (-3,4)

OpenStudy (anonymous):

ty, i have another, but its more confusing, you think u could help me with it

OpenStudy (anonymous):

Find the maximum or minimum of the following quadratic function: y = 16x^2+ 40x + 25. A. 0 B. 25 C. -25 D. 50

OpenStudy (anonymous):

@texaschic101